# MSHyper: Multi-Scale Hypergraph Transformer for Long-Range Time Series Forecasting

### v1

@@ 5.3 Ablation Studies
0.405
0.427
0.428
0.439
0.503
0.507
5.4 Parameter Studies
We also perform parameter studies to measure the impact of the aggregation function, $k$-hop, and input length on ETTh1 dataset. The experimental results are shown in Table 5, Figure 5, and Figure 5, from which we can observe that: 1) The results based on the convolution operation outperform those based on other operations, possibly because convolution kernels can extract more complex features. 2) The best performance can be obtained when the $k$-hop is 4. The reason may be that a small $k$-hop limits interactions to the neighboring nodes, and a large $k$-hop would introduce noises. 3) MSHyper achieves the optimal performance when the input length is 168. The reason may be that the sequence with a shorter input length contains insufficient pattern information, while the sequence with a longer input length introduces more noise interference for high-order pattern interactions.
Table 5: Impact of aggregation function.
Methods Prediction length
168 336 720
Max-pooling
@@ 5.4 Parameter Studies
0.508
0.468
0.437
0.457
0.513
0.524
Average-pooling
@@
0.532
0.506
0.431
0.455
0.512
0.523
Conv
MSE
MAE
0.405
0.427
0.428
0.439
0.503
0.507
Figure 4: Impact of $k$-hop.
Refer to caption k-hop
MSE
Refer to caption Input length
MSE
Figure 4: Impact of $k$-hop.
Figure 5: Impact of input length.
6 Conclusions
In this paper, we propose MSHyper for long-range time series forecasting. Specifically, the H-HGC module is introduced to provide foundations for modeling high-order interactions between temporal patterns. The TMP mechanism is employed to aggregate high-order pattern information and learn the interaction strength between temporal patterns of different scales. Extensive experiments on five real-world datasets show the superiority of MSHyper.

### v2

@@ V-D Ablation Studies
0.383
0.392
0.435
0.423
0.480
0.445
0.482
0.467
V-E Case Studies
To evaluate the prediction performance of different models, we plot the forecasting results on Weather and Electricity datasets for qualitative comparison. The visualization of forecasting results are shown in Figure 4 and 5, from which we can observe that the existing models have poorer fitting ability on Weather dataset than on Electricity dataset. To explore the reason, Figure 6 shows the auto-correlation graphs of sampled variables on Electricity and Weather datasets. On Electricity dataset, clear daily and weekly patterns can be observed, while on Weather dataset, it is difficult to identify obvious daily or weekly patterns. In addition, MSHyper is still able to achieve good predictive performance on Weather dataset, which may be due to its ability to capture more diverse long-term and short-term patterns by modeling high-order temporal pattern interactions.
(a) DLinear
(b) TimesNet
(c) MSGNet
(d) MSHyper
Figure 4: Forecasting results of different models on Electricity dataset under the input-96-predict-96 setting. The black line represents the ground truth and the orange line represents the predicted results.
(a) DLinear
(b) TimesNet
(c) MSGNet
(d) MSHyper
Figure 5: Forecasting results of different models on Weather dataset under the input-96-predict-96 setting. The black line represents the ground truth and the orange line represents the predicted results.
(a) Electricity dataset
(b) Weather dataset
Figure 6: The auto-correlation graphs on Electricity and Weather datasets.
V-F Parameter Studies
Input Length. To investigate the impact of input length, we set input length $T\in\{48,96,192,336\}$ and record the results of MSHyper for multivariate long-range time series forecasting. The experimental results are shown in Figure 7, from which we can observe that with the increase of input length, the performance of MSHyper shows an upward trend. The reason is that MSHyper can capture more historical pattern information with the increase of input length.
Input length
@@ V-F Parameter Studies
Figure 7: Results of MSHyper with different input length on ETTm1 dataset.
$\boldsymbol{k}$-hop. We also perform parameter studies to measure the impact of the $k$-hop on ETTm1 dataset. Figure 8 shows the results of MSHyper on ETTm1 dataset by varying the $k$-hop from 2 to 5, form which we can observe that the best performance can be obtained when the $k$-hop is 3. The reason is that a small $k$-hop limits interactions to the neighboring nodes, and a large $k$-hop would introduce noises.
k-hop
@@
Figure 8: Results of MSHyper with different $k$-hop on ETTm1 dataset.
V-G Computation Cost
To evaluate the computation cost of MSHyper, we compare the training time, GPU occupation, and forecasting performance of MSHyper, MSGNet, TimesNet, Crossformer, and DLinear on Electricity dataset with the forecasting horizon of 96. The experimental results are shown in Table VIII, from which we can observe that DLinear has lower GPU occupation and runs fastest in these methods. But it gets worse forecasting performance than MSHyper, MSGNet, and TimesNet. Compared with MSGNet, TimesNet, and Crossformer, MSHyper runs faster and gets the best forecasting performance. Overall, comprehensively considering the significant forecasting performance improvement and the computation cost, MSHyper demonstrates the superiority over existing methods.
TABLE VIII: Computation costs of different methods.
Methods Training Time/epoch GPU Occupation MSE MAE
MSHyper 13.51s 2186 0.152 0.252
MSGNet 1519.10s 11550 0.165 0.274
TimesNet 587.35s 3726 0.168 0.272
Crossformer 88.36s 9468 0.219 0.314
DLinear 4.35s 756 0.197 0.282
VI Conclusions and Future Work
In this paper, we propose MSHyper for long-range time series forecasting. Specifically, the H-HGC module is introduced to provide foundations for modeling high-order interactions between temporal patterns. The TMP mechanism is employed to aggregate high-order pattern information and learn the interaction strength between temporal patterns of different scales. Extensive experiments on eight real-world datasets show the superiority of MSHyper.
In future work, it is of interest to extend MSHyper in the following two aspects: First, since the prior knowledge-based hypergraph structures may limit the scalability of MSHyper, we will design an adaptive hypergraph module to learn the best hypergraph structures. Second, to tackle complex scenariors with diverse temporal pattern interactions, we will design a neural architecture search framework to search the best combination of hypergraph structures, capturing high-order correlations among temporal patterns of different scales automatically.